OpenStudy (anonymous):

help me guys! find the equation of the tangent line at (2,-1) on the circle (x+20)^2+(y+2)^2=25

OpenStudy (anonymous):

hint : tangent line is perpendicular to the radial line..

OpenStudy (anonymous):

y=2x-5 working tangent is perpendicular to the radius of the circle. Gradient at the point (2,-1) from origin (0,0) = -1/2 rise/run or (y1-y2)/(x1-x2) Since it is perpendicular the gradient of the tangent is equal to the negative reciprocal of the gradient of the circle. Gradient of tangent = 2 so equation y=2x+c Find c by substituting the values (2,-1) -1= 2(2)+c c=-5